OpenStudy (gokuporter):

Write the equation of the line in standard form. Points on the line: (-5,2) and (1/2,4)

OpenStudy (gokuporter):

@jagr2713

OpenStudy (trojanpoem):

The standard form of line equation has the following formula: Ax + By = C y = mx + b ( Slope -intercept form) m = dy/dx = 4 - 2/ (0.5) - (-5) = 4/11 y = 4/11 x + b 2 = 4/11 * -5 + b b = 42/11 y = 4/11 x + 42/11 11y = 4x + 42 4x - 11y = -42

OpenStudy (trojanpoem):

By the way, y = -A/B x + C/B so slope = -A/B , y-intercept = C/B You can make 3 equations -A/B , 2 points and form the eqn the same way.

OpenStudy (gokuporter):

Okay, so the Point Intercept form is y = 4/11 x +b. How did you get 4/11 for the slope and b = 42/11.

OpenStudy (gokuporter):

@TrojanPoem

OpenStudy (trojanpoem):

I used the slope -intercept form. After finding the slope, just plug in the ordered pair. and solve for b.

OpenStudy (gokuporter):

How did you get for 4/11 for the slope?

OpenStudy (trojanpoem):

Points on the line: (-5,2) and (1/2,4) \[\frac{ \Delta y }{ \Delta x } = \frac{ 4 - 2 }{ 0.5 - (-5) } = \frac{ 4 }{ 11 }\]

OpenStudy (gokuporter):

0.5=1/2?

OpenStudy (trojanpoem):

Yeah.

OpenStudy (gokuporter):

4/11 * -5 = 42/11?

OpenStudy (trojanpoem):

2 - (4/11 * -5) = 42/11

OpenStudy (gokuporter):

Is 4/11 * 5 20/11?

OpenStudy (trojanpoem):

Yeah, as I assume that 4/11 * 5 = (4/11) *5

OpenStudy (trojanpoem):

\[\frac{ 4 }{ 11 } * 5\]

OpenStudy (gokuporter):

2 - 20/11 .-.

OpenStudy (trojanpoem):

2/11 , (assuming you're not referring to your question) 2 - (4/11) * -5 = 2 + 20/11 = 22/11 + 20/11 = 42/11

OpenStudy (gokuporter):

Thanks for sorting that out. :)

OpenStudy (trojanpoem):

Never mind.

OpenStudy (gokuporter):

Why'd you put 11y?

OpenStudy (trojanpoem):

You have to put it as Ax + By = C maybe - 4//1x + y = 42/11 ( Is ok! )